Circular No. 6882

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998bm IN IC 2458 W. D. Li, M. Modjaz, R. R. Treffers, and A. V. Filippenko, University

of California at Berkeley (UCB), report their discovery of an apparent supernova during the course of the Lick Observatory Supernova Search (cf. IAUC 6627) with the 0.8-m Katzman Automatic Imaging Telescope. SN 1998bm was found and confirmed on unfiltered CCD images obtained on Apr. 21.2 (mag ∼ 17.6) and 22.2 UT (mag ∼ 17.4), respectively. The object is located at α = 9^{h}21^{m}30^{s}.49, δ = +64^{o}14′25′′.4 (equinox 2000.0), which is ∼ 1′′ east and 5′′ north of the nucleus of IC 2458. An image of the same field taken on 1996 Nov. 13 (limiting mag ∼ 18.5) by the Beijing Astronomical Observatory Supernova Survey shows no star at this position.

A. V. Filippenko and D. C. Leonard, UCB, report that a spectrum obtained on Apr. 23 with the Shane 3-m reflector at Lick Observatory shows that SN 1998bm is of type II; the hydrogen Balmer lines have P-Cyg profiles.

XTE J1906+09 D. Marsden and R. Rothschild, Center for Astrophysics and Space Sci-

ences, University of California at San Diego, report for the RXTE/HEXTE team: “We report the discovery of a pulsating x-ray source designated XTE J1906+09 during observations with RXTE of a 1^{o} (FWHM) region on 1996 Aug. 16–19 containing the soft γ-ray repeater SGR 1900+14. The RXTE field-of-view was centered at α = 19^{h}05^{m}43^{s}, δ = +8^{o}58′.8 (equinox 2000.0). Coherent pulsations were detected from the source by both the PCA and HEXTE instruments at a fundamental period of 89.17 ± 0.02 s, with higherorder harmonics visible in the 2–10-keV power spectrum. The folded pulsar lightcurve is energy dependent, with a multipeaked morphology (2–20 keV) at lower energies that becomes singly peaked in the energy range 20–200 keV. After correcting for instrumental, sky, and Galactic ridge backgrounds, the pulsed fractions are 16, 6, and 4 percent in the energy ranges 2–10, 10– 20, and 20–300 keV, respectively. The phase-averaged spectrum is well fitted over the entire energy range by either a power law of photon index 1.9 ± 0.1 or a thermal bremsstrahlung model of temperature kT = 20 ± 5 keV, with a 2–10-keV flux of ∼ 1.5 × 10^{−11} erg cm^{−2} s^{−1}. The column density implied by the spectral fits is N_{H} ∼ 1 × 10^{23} cm^{−2}, implying a distance to the source of ∼ 10–20 kpc. We encourage observations at other wavelengths to refine the position and search for a counterpart.”

1998 April 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green